# Quaternionic Pole Placement via Companion Forms and the Ackermann Formula

### v1

@@ Abstract
We present an extension of state-feedback pole placement for quaternionic systems, based on companion forms and the Ackermann formula. For controllable single-input quaternionic LTI models, we define a companion polynomial that right-annihilates its companion matrix, characterize spectra via right-eigenvalue similarity classes, and prove coefficient-matching design in controllable coordinates. We then derive a coordinate-free Ackermann gain expression valid for real target polynomials, and state its scope and limitations. Short examples demonstrate correctness, practical use, and numerical simplicity.
@@ I Introduction
These workflows resemble their $\mathbb{R}/\mathbb{C}$ counterparts—companion form, one-shot gain, and an Ackermann-style formula—but they differ in substance: over $\mathbb{H}$, right eigenvalues (the poles) are similarity classes (not points), polynomial evaluation is side/ordering-sensitive, and the Ackermann step is valid only for real (central) coefficients. Our design preserves the familiar flow while making these quaternionic nuances explicit.
The note is organized as follows. Section II reviews quaternion algebra and matrix preliminaries; Section III states the quaternionic LTI model. Section IV introduces the controllable companion form and its companion polynomial. Section V formulates pole placement over $\mathbb{H}$. Section VI presents coefficient matching in controllable coordinates; Section VII derives the quaternionic Ackermann formula and delineates its scope; and Section VIII concludes.
@@ II Quaternions and Matrices
Two quaternions $q,r$ are similar, written $q\sim r,$ if $q=\alpha^{-1}r\alpha$ for some nonzero $\alpha\in\mathbb{H}.$ Similarity preserves $\operatorname{Re}(\cdot),$ $\left|\cdot\right|,$ and $\left|\operatorname{Im}(\cdot)\right|$ and partitions $\mathbb{H}$ into equivalence classes $\left[\cdot\right]$: real quaternions are isolated classes $\left[a\right]=\{a\}$; If $\operatorname{Im}(q)\neq 0,$ then $[q]$ is a 2-sphere $\{\mathrm{Re}(q)+u\,\left|\operatorname{Im}(q)\right|:u\in\mathrm{span}\{\mathbf{i},\mathbf{j},\mathbf{k}\},\left|u\right|=1\}$ that meets the standard complex line $\{a+b\mathbf{i}:\,a,b\in\mathbb{R}\}$ in exactly two points $\operatorname{Re}(q)\pm i\,\left|\operatorname{Im}(q)\right|.$ For example, $\mathbf{k}]$ contains all imaginary units, including $\pm\mathbf{i}.$
@@
Let $\mathbb{H}[\lambda]$ denote the ring of quaternionic polynomials $p(\lambda)=\sum_{k=0}^{n}p_{k}\lambda^{k}$ with $p_{k}\in\mathbb{H}$, where the indeterminate $\lambda$ commutes with all coefficients. Evaluation at a quaternion $q\in\mathbb{H}$, however, is order–dependent: $p(q)_{\mathscr{r}}:=\sum_{k=0}^{n}p_{k}q^{k}$ is the right value and $p(q)_{\ell}:=\sum_{k=0}^{n}q^{k}p_{k}$ is the left value.
@@
Isolated and spherical zeros
A quaternionic polynomial has either $n$ (not necessarily distinct) right zeros or infinitely many. A nonreal zero is either isolated (a single point of its class) or spherical (the entire class $[q]$, a 2-sphere). The spherical case occurs exactly when the real quadratic $\chi_{q}(\lambda)=\lambda^{2}-2\,\mathrm{Re}(q)\,\lambda+|q|^{2}$ divides $p(\lambda);$ Real zeros are always isolated since $r]=\{r\}.$ With real coefficients, all zeros are either real (hence isolated) or spherical [20].
@@
We view column vectors as a right $\mathbb{H}$-module. Let $e_{i}$ denote the $i$th canonical basis vector in $\mathbb{R}^{n}\subset\mathbb{H}^{n}$, i.e., the column with a $1$ in position $i$ and zeros elsewhere; in particular $e_{n}=[0,\ldots,0,1]^{\top}$. Since these entries are real (central), right/left multiplication with scalars is unambiguous when $e_{i}$ appears.
@@ IV Companion Form and Companion Polynomial
Let $A\in\mathbb{H}^{n\times n}$ and $B\in\mathbb{H}^{n}$ be given, and assume $\mathcal{C}=[B,AB,\ldots,A^{n-1}B]$ is invertible over $\mathbb{H}$.
@@
    $A_{c}=T^{-1}AT\quad\textrm{ and }\quad B_{c}$ $=$ $T^{-1}B$ (6)
    and the pair $(A_{c},B_{c})$ is in lower controllable companion form (4), where the coefficients $a_{0},\dots,a_{n-1}\in\mathbb{H}$ are uniquely defined by the (right) expansion
    $A^{n}B\;=\;-\sum_{\mathbf{k}=0}^{n-1}A^{k}B\,a_{k}.$ (7)
@@
    $T^{-1}$ $=$ $\begin{bmatrix}t\\ tA\\ \vdots\\ tA^{n-1}\end{bmatrix}$. (8)
@@
The proof adapts the standard companion argument for real matrices (see [2]) to the quaternionic setting.
First, note that the coefficients $a_{i}$ defined by (7) exist and are unique whenever $\mathcal{C}$ is invertible.
Now with (9),(8) and the definition of $\mathcal{C}$ we have
$\big(T^{-1}\mathcal{C}\big)_{ij}=tA^{i+j-2}B.$ (10)
Since
[…]
it follows that the matrix $T^{-1}\mathcal{C}$ is unit anti-triangular (ones on the anti-diagonal, zeros above).
The claim that $(T^{-1}AT,\;T^{-1}B)$ is in lower controllable canonical form — and hence equals $(A_{c},B_{c})$ — is proved in three steps:
Step 1 - Input vector: The first column of $\mathcal{C}$ is $B,$ hence
$T^{-1}B$ $=$ $\begin{bmatrix}tB\\ tAB\\ \vdots\\ tA^{n-1}B\end{bmatrix}=\begin{bmatrix}0\\ \vdots\\ 0\\ 1\end{bmatrix}=e_{n}=B_{c}.$
Step 2 - Superdiagonal ones: Denoting
$M$ $\coloneqq$ $T^{-1}AT,$
we prove $M$ has ones above the main diagonal.
Take the reversed controllability basis
$W$ $=$ $[A^{n-1}B,\;A^{n-2}B,\;\dots,\;B],$
which is invertible and set
$Q$ $\coloneqq$ $T^{-1}W,$
which is also invertible.
For $i=1,\dots,n-1$ and any $j=1,\dots,n$, the $(i,j)$ element of the matrix $MQ$ reads
$e_{i}^{\mathsf{T}}M\,Qe_{j}$ $=$ $e_{i}^{\mathsf{T}}T^{-1}ATT^{-1}We_{j}=e_{i}^{\mathsf{T}}T^{-1}AWe_{j}$
$=$ $e_{i}^{\mathsf{T}}T^{-1}A^{n-j+1}B=tA^{n+i-j}B,$
since $e_{i}^{\mathsf{T}}T^{-1}=tA^{i-1}$ and $AWe_{j}=A^{n-j+1}B$.
Similarly, the $(i+1,j)$ element of the matrix $Q$ reads
$e_{i+1}^{\mathsf{T}}Qe_{j}$ $=$ $e_{i+1}^{\mathsf{T}}T^{-1}We_{j}=tA^{n+i-j}B,$
since $e_{i+1}^{\mathsf{T}}T^{-1}=tA^{i}$ and $We_{j}=A^{n-j}B$.
Hence
$(e_{i}^{\mathsf{T}}M-e_{i+1}^{\mathsf{T}})\,Qe_{j}=0\textrm{ for all }j.$
Because $Q$ is invertible, we conclude $e_{i}^{\top}M=e_{i+1}^{\top}$ for $i=1,\ldots,n-1$; i.e., $M$ has ones on the superdiagonal and zeros above.
Step 3 - Last row (shift-register argument): Apply $T^{-1}$ to (7) and use $M^{k}e_{n}=T^{-1}A^{k}B$ (where $e_{n}=T^{-1}B$ commutes with $a_{k}$):
$M^{n}e_{n}\;=\;-\sum_{k=0}^{n-1}M^{k}e_{n}\,a_{k}.$ (12)
Set $v_{k}:=M^{k}e_{n}$. Since $M$ is a shift matrix, the first n1n\!-\!1 rows only shift upward; thus for each row $i$ there is a unique k=nik=n-i with $(v_{n-i})_{i}=1$ and $(v_{k})_{i}=0$ for $k\neq n-i$. Reading (12) in row $i$ gives $(v_{n})_{i}=-(v_{n-i})_{i}a_{n-i}=-a_{n-i}.$ But $(v_{n})_{i}\!=\!r_{i-1}$ (it is created by the last row via the shifts). Hence $r_{i-1}\!=\!-a_{n-i}$ for all $i$, i.e., the last row of $M$ is $a_{0},\,-a_{1},\,\dots,\,-a_{n-1}\,]$.
Therefore $M$ is the lower controllable companion form and $A_{c}:=M=T^{-1}AT,\ B_{c}:=T^{-1}B=e_{n}$.
Step 4 - Uniqueness: Since $\mathcal{C}$ is invertible, the expansion $A^{n}B=-\sum_{k=0}^{n-1}A^{k}B\,a_{k}$ determines $a_{0},\dots,a_{n-1}$ uniquely, hence $A_{c}$ is unique. With $B_{c}=e_{n}$ and ones on the superdiagonal fixed, the companion controllability matrix $\mathcal{C}_{c}=[B_{c},A_{c}B_{c},\dots,A_{c}^{n-1}B_{c}]$ is uniquely determined, and $T^{-1}\mathcal{C}=\mathcal{C}_{c}$ implies $T^{-1}=\mathcal{C}_{c}\,\mathcal{C}^{-1},$ so $T$ is unique as well. ∎
@@
annihilates $A_{c}$ from the right in the sense that
@@
Set $v_{k}:=A_{c}^{\,k}e_{n}$, $k\geq 0$ and collect the first $n$ of them into the Krylov matrix
$\mathcal{K}:=[\,v_{0},\;v_{1},\;\dots,\;v_{n-1}\,]\in\mathbb{H}^{n\times n}.$
For the companion $A_{c}$, the columns $v_{0},\dots,v_{n-1}$ are linearly independent (indeed $\mathcal{K}$ is a permutation of the identity plus lower-row corrections), hence $\mathcal{K}$ is invertible.
By construction of $A_{c}$,
$A_{c}v_{k}$ $=$ $v_{k+1}\qquad(0\leq k\leq n-2),$
$A_{c}v_{n-1}$ $=$ $v_{n}=-\sum_{k=0}^{n-1}a_{k}\,v_{k},$ (21)
where the last equality is exactly the “feedback row” $a_{0},\dots,-a_{n-1}]$ applied at the wrap-around.
Left-multiplying the second relation in (21) by $A_{c}$ gives, for all $j\geq 0$
$v_{n+j}=A_{c}^{\,j}v_{n}=-\sum_{k=0}^{n-1}a_{k}\,A_{c}^{\,j}v_{k}=-\sum_{k=0}^{n-1}a_{k}\,v_{k+j}.$ (22)
Now evaluate the polynomial $a(\lambda)$ from the right on $A_{c}$ and apply it to each $v_{j}$ with $0\leq j\leq n-1$:
$\bigl(A_{c}^{n}+a_{n-1}A_{c}^{n-1}+\cdots+a_{0}I\bigr)v_{j}=v_{n+j}+\sum_{k=0}^{n-1}a_{k}v_{k+j}=0$
by (22). Thus every column of $a(A_{c})\mathcal{K}$ is zero, i.e. $a(A_{c})\mathcal{K}=0.$ Since $\mathcal{K}$ is invertible, $a(A_{c})=0$, which is the claim. ∎
@@
The right spectrum of the companion matrix consists of similarity classes formed by the right zeros of the companion polynomial.
@@
    Consequently, $\sigma_{\mathscr{r}}(A_{c})=\bigcup_{i}[\lambda_{i}]$, where $\{\lambda_{i}\}$ is the multiset of (not necessarily distinct) right zeros of $a$.
@@
Step 1: Put $v\!=\!\left[1,\lambda,\ldots,\lambda^{n-1}\right]^{\mathsf{T}}$ and assume $a(\lambda)\!=\!0$. By construction,
@@
Step 2: By Step (1), every right zero $\lambda_{i}$ yields a right eigenpair $A_{c}v=v\lambda_{i}$. Hence the whole similarity class $\lambda_{i}]$ lies in the right spectrum $\lambda_{i}]\subset\sigma_{\mathscr{r}}(A_{c}).$
For the converse, take a right eigenpair $A_{c}x=x\mu,$ $x\neq 0.$ From the first n1n-1 rows,
$x_{1}$ $=$ $x_{0}\mu,\ x_{2}=x_{1}\mu=x_{0}\mu^{2},\ \ldots,\ x_{n-1}=x_{0}\mu^{n-1}.$
As above $x_{0}\neq 0$. The last row gives
$-(a_{0}x_{0}+a_{1}x_{0}\mu+\cdots+a_{n-1}x_{0}\mu^{n-1})$ $=$ $x_{0}\mu^{n}.$
Right-multiply by $x_{0}^{-1}$
$-a_{0}-a_{1}(x_{0}\mu x_{0}^{-1})-\cdots-a_{n-1}(x_{0}\mu^{n-1}x_{0}^{-1})$ $=$ $x_{0}\mu^{n}x_{0}^{-1}.$
Let $\lambda:=x_{0}\mu x_{0}^{-1}.$ Then $x_{0}\mu^{k}x_{0}^{-1}=\lambda^{k}$ for all $k$, so the equality becomes
$-a_{0}-a_{1}\lambda-\cdots-a_{n-1}\lambda^{n-1}=\lambda^{n}\;\iff\;a(\lambda)=0.$
Thus every right eigenvalue $\mu$ is similar to a right zero $\lambda$ of $a(\lambda)$, i.e. $\mu\in[\lambda]$. Hence
$\sigma_{\mathscr{r}}(A_{c})$ $=$ $\textstyle{\bigcup}_{i}\,[\lambda_{i}],$
with $\{\lambda_{i}\}$ the multiset of right zeros of $a(\lambda).$
Step 3: The left zeros of $a(\lambda)$ are known to be similar to the right zeros of $a(\lambda)$ [9] and as such are also in the right spectrum of $A_{c}.$
@@
Equivalently, the coefficients $a_{0},\dots,a_{n-1}\in\mathbb{H}$ are the unique solution of
$A^{n}B\;=\;-\sum_{k=0}^{n-1}A^{k}B\,a_{k},$
which is guaranteed by the invertibility of $\mathcal{C}$.
@@
Throughout this note, we adopt the lower controllable companion form. Other orientations—upper, right, or left—are equivalent; using any of them leads to the same results (up to the obvious permutation/similarity). In particular, the companion polynomial is independent of the chosen orientation: all companion forms yield the same monic polynomial.
@@
    Consequently, $\sigma_{\mathscr{r}}(A)=\bigcup_{i}[\lambda_{i}]$, where $\{\lambda_{i}\}$ is the multiset of (not necessarily distinct) right zeros of $a$.
@@
Hence, the companion polynomial shares some properties of the minimal/characteristic polynomial even in quaternions, but not all! In particular, it does not annihilate $A$ in general; it only annihilates its companion form $A_{c}$. Hence $a(A_{c})=0$ but, in general, $a(A)\neq 0.$
@@
Having associated a companion polynomial to a controllable pair $(A,B)$, we may—when $A$ is cyclic—also speak of a companion polynomial of $A$. This polynomial is not intrinsic to $A$: different cyclic choices of $B$ generally produce different polynomials. Nevertheless, their right zeros agree at the level of quaternionic similarity classes: for every such $B$, the right zeros of the associated companion polynomial form the same multiset of conjugacy classes, namely the right spectrum of $A$. By contrast, once a specific pair $(A,B)$ is fixed, the lower companion realization is unique, and therefore the polynomial attached to that pair—equivalently, to its $A_{c}$—is unique.

### v2

@@ Abstract
We present an extension of state-feedback pole placement for quaternionic systems, based on companion forms and the Ackermann formula. For controllable single-input quaternionic LTI models, we define a companion polynomial that annihilates its companion matrix, characterize spectra via right-eigenvalue similarity classes, and prove coefficient-matching design in controllable coordinates. We then derive a coordinate-free Ackermann gain expression valid for real target polynomials, and state its scope and limitations. Short examples demonstrate correctness, practical use, and numerical simplicity.
@@ I Introduction
These workflows mirror their $\mathbb{R}/\mathbb{C}$ counterparts—companion coordinates, coefficient matching, and an Ackermann-type gain—yet key quaternionic differences remain: right poles are similarity classes, polynomial evaluation is order-sensitive, and the Ackermann step requires real (central) target coefficients. We retain the familiar design flow while making these nuances explicit.
The note is organized as follows. Sections II–IV develop the quaternionic preliminaries, LTI model, and companion-form framework. Sections V–VII treat pole placement, coefficient matching, and the Ackermann formula, respectively; Section VIII concludes.
@@ II Quaternions and Matrices
Two quaternions $q,r$ are similar, written $q\sim r,$ if $q=\alpha^{-1}r\alpha$ for some nonzero $\alpha\in\mathbb{H}.$ Similarity preserves $\operatorname{Re}(\cdot),$ $\left|\cdot\right|,$ and $\left|\operatorname{Im}(\cdot)\right|$ and partitions $\mathbb{H}$ into equivalence classes $\left[\cdot\right]$: real quaternions are isolated classes $\left[a\right]=\{a\}$; if $\operatorname{Im}(q)\neq 0,$ then $[q]$ is a 2-sphere $\{\mathrm{Re}(q)+u\,\left|\operatorname{Im}(q)\right|:u\in\mathrm{Im}(\mathbb{H})\cong\mathbb{R}^{3},\ |u|=1\}$ that meets the standard complex line $\{a+b\mathbf{i}:\,a,b\in\mathbb{R}\}$ in exactly two points $\operatorname{Re}(q)\pm i\,\left|\operatorname{Im}(q)\right|.$ For example, $\mathbf{k}]$ contains all imaginary units, including $\pm\mathbf{i}.$
@@
Let $\mathbb{H}[\lambda]$ denote the ring of quaternionic polynomials $p(\lambda)=\sum_{k=0}^{n}p_{k}\lambda^{k}$ with $p_{k}\in\mathbb{H}$, where the indeterminate $\lambda$ commutes with all coefficients. Evaluation is order–dependent: for $q\in\mathbb{H}$ and $A\in\mathbb{H}^{m\times m}$ we define
$p(q)_{\mathscr{r}}$ $=\!\sum_{k=0}^{n}p_{k}q^{k},\quad p(q)_{\ell}\!=\!\sum_{k=0}^{n}q^{k}p_{k},$
$p(A)_{\mathscr{r}}$ $=\!\sum_{k=0}^{n}p_{k}A^{k}.$
i.e., in the right matrix evaluation, the matrix powers appear on the right of the coefficients. Analogously one may set $p(A)_{\ell}:=\sum_{k=0}^{n}A^{k}p_{k}$ if needed. If $p\in\mathbb{R}[\lambda]$ (central coefficients), then $p(\cdot)_{\mathscr{r}}=p(\cdot)_{\ell}$ and we write simply $p(\cdot)$; otherwise we keep the evaluation index explicit.
@@
Isolated and spherical right zeros
Let $p\in\mathbb{H}[\lambda]$ be nonzero with $n=\deg p$. A nonreal right zero is either isolated (a single point of its similarity class) or spherical (the whole class $[q]$, hence infinitely many zeros). The spherical case occurs exactly when the real quadratic $\chi_{q}(\lambda)=\lambda^{2}-2\operatorname{Re}(q)\lambda+|q|^{2}\in\mathbb{R}[\lambda]$ divides $p(\lambda)$; thus $p$ has infinitely many right zeros iff it has a spherical right zero. If $p$ has no spherical right zeros, then all right zeros are isolated and their (algebraic) multiplicities sum to $n=\deg p$. In general, the same total $n$ is obtained by counting isolated zeros with (algebraic) multiplicity and counting each spherical class twice (with spherical multiplicity). [17, 14]
@@
We view $\mathbb{H}^{n}$ as a right $\mathbb{H}$-module. Let $e_{i}\in\mathbb{R}^{n}\subset\mathbb{H}^{n}$ be the $i$th canonical basis vector (so $e_{n}=[0,\ldots,0,1]^{\top}$). Since $e_{i}$ has real entries, left/right scalar multiplication is unambiguous.
@@ IV Companion Form and Companion Polynomial
Let $A\in\mathbb{H}^{n\times n}$ and $B\in\mathbb{H}^{n}$ be given, and assume the controllability matrix $\mathcal{C}=[B,AB,\dots,A^{n-1}B]$ is invertible over $\mathbb{H}$.
@@
    $A_{c}=T^{-1}AT,\qquad B_{c}=T^{-1}B,$ (6)
    and the pair $(A_{c},B_{c})$ is in lower controllable companion form (4). In particular, the coefficients $a_{0},\dots,a_{n-1}\in\mathbb{H}$ are defined by the last row of $A_{c}$,
    $e_{n}^{\top}A_{c}=\begin{bmatrix}-a_{0}&-a_{1}&\cdots&-a_{n-1}\end{bmatrix}$.
@@
    $T^{-1}$ $=$ $\begin{bmatrix}t\\ tA\\ \vdots\\ tA^{n-1}\end{bmatrix}$, (7)
@@
The proof follows the standard companion-form argument [2], adapted to $\mathbb{H}$.
Since $\mathcal{C}$ is invertible, $t=e_{n}^{\top}\mathcal{C}^{-1}$ is well defined, and so is $T^{-1}$ from (7). For $i,j=1,\dots,n$,
$(T^{-1}\mathcal{C})_{ij}=tA^{i+j-2}B.$
Because $t$ is the last row of $\mathcal{C}^{-1}$, we have $tA^{k}B=0$ for $k=0,\dots,n-2$ and $tA^{n-1}B=1$. Hence $T^{-1}\mathcal{C}$ is unit anti-triangular, so $T$ is invertible.
Set $M:=T^{-1}AT$. First, since the first column of $\mathcal{C}$ is $B$,
$T^{-1}B=\begin{bmatrix}tB&tAB&\cdots&tA^{n-1}B\end{bmatrix}^{\!\top}=e_{n}.$
Next, let $W=[A^{n-1}B,\dots,B]$ and $Q:=T^{-1}W$. Since both $T$ and $W$ are invertible, so is $Q$. For $i=1,\dots,n-1$ and $j=1,\dots,n$,
$e_{i}^{\top}MQe_{j}=e_{i}^{\top}T^{-1}AWe_{j}=tA^{n+i-j}B,$
while
$e_{i+1}^{\top}Qe_{j}=e_{i+1}^{\top}T^{-1}We_{j}=tA^{n+i-j}B.$
Thus $(e_{i}^{\top}M-e_{i+1}^{\top})Qe_{j}=0$ for all $j$, and invertibility of $Q$ gives $e_{i}^{\top}M=e_{i+1}^{\top}$ for $i=1,\dots,n-1$. Therefore the first n1n-1 rows of $M$ are exactly the shift rows of the lower controllable companion form.
Writing the last row as $e_{n}^{\top}M=[-a_{0},\,-a_{1},\,\dots,\,-a_{n-1}]$, we conclude that $M$ has the form (4). Hence $A_{c}=T^{-1}AT$ and $B_{c}=T^{-1}B=e_{n}$.
For uniqueness, let $\widetilde{T}$ be any invertible matrix such that $\widetilde{T}^{-1}B=e_{n}$ and $\widetilde{M}:=\widetilde{T}^{-1}A\widetilde{T}$ is in lower controllable companion form. Put $\widetilde{t}:=e_{1}^{\top}\widetilde{T}^{-1}$. Since $e_{i}^{\top}\widetilde{M}=e_{i+1}^{\top}$ for $i=1,\dots,n-1$, we get $e_{i}^{\top}\widetilde{T}^{-1}=\widetilde{t}A^{i-1}$, so $\widetilde{T}^{-1}$ has the same form as (7), with $\widetilde{t}$ in place of $t$.
Now
$\widetilde{t}\,\mathcal{C}=e_{1}^{\top}\widetilde{T}^{-1}\mathcal{C}=\begin{bmatrix}e_{1}^{\top}e_{n}&e_{1}^{\top}\widetilde{M}e_{n}&\cdots&e_{1}^{\top}\widetilde{M}^{n-1}e_{n}\end{bmatrix}$.
Because $\widetilde{M}$ has the companion shift rows, $e_{1}^{\top}\widetilde{M}^{k}=e_{k+1}^{\top}$ for $k=0,\dots,n-1$, hence $\widetilde{t}\,\mathcal{C}=e_{n}^{\top}$. Therefore $\widetilde{t}=e_{n}^{\top}\mathcal{C}^{-1}=t$, so $\widetilde{T}^{-1}=T^{-1}$, $\widetilde{T}=T$, and the pair $(A_{c},B_{c})$ is unique. ∎
@@
satisfies the right matrix evaluation identity
@@
Define the row sequence
$r_{k}$ $:=$ $e_{1}^{\top}A_{c}^{k},\qquad k\geq 0.$
Then
$r_{0}=e_{1}^{\top},\;r_{1}=e_{2}^{\top},\;\ldots,\;r_{n-1}=e_{n}^{\top},$
so $r_{0},\dots,r_{n-1}$ are exactly the rows of the identity matrix and in particular form a basis of $\mathbb{H}^{1\times n}$.
By the shape of the companion matrix,
$r_{n}$ $=$ $e_{n}^{\top}A_{c}=\big[-a_{0},\dots,-a_{n-1}\big]=-\sum_{k=0}^{n-1}a_{k}\,r_{k}.$
Since $r_{k+1}=r_{k}A_{c}$, it follows that for every $j\geq 0$,
$r_{n+j}$ $=$ $r_{n}A_{c}^{j}=-\sum_{k=0}^{n-1}a_{k}\,r_{k}A_{c}^{j}=-\sum_{k=0}^{n-1}a_{k}\,r_{k+j}.$ (16)
Now evaluate $a(\lambda)$ at $A_{c}$ using the right matrix evaluation:
$a(A_{c})_{\mathscr{r}}$ $=$ $a_{0}I+a_{1}A_{c}+\cdots+a_{n-1}A_{c}^{n-1}+A_{c}^{n}.$
For each $j=0,\dots,n-1$,
$r_{j}a(A_{c})_{\mathscr{r}}$ $=$ $a_{0}r_{j}+a_{1}r_{j+1}+\cdots+a_{n-1}r_{j+n-1}+r_{j+n}$
$=$ $0$
by (16). Hence every row $r_{j}$, $j=0,\dots,n-1$, annihilates $a(A_{c})_{\mathscr{r}}$. Since $r_{0},\dots,r_{n-1}$ are the rows of the identity matrix, this implies $a(A_{c})_{\mathscr{r}}=0$. ∎
@@
The right spectrum of the companion matrix is the set of similarity classes determined by the right zeros of the companion polynomial.
@@
    Consequently,
    $\sigma_{\mathscr{r}}(A_{c})=\{[\lambda]:a(\lambda)_{\mathscr{r}}=0\}.$
    Equivalently, if $\lambda_{1},\dots,\lambda_{m}$ are the distinct right zeros of $a$ modulo similarity, then
    $\sigma_{\mathscr{r}}(A_{c})=\{[\lambda_{1}],\dots,[\lambda_{m}]\}.$
@@
Step 1: Put $v=\left[1,\lambda,\ldots,\lambda^{n-1}\right]^{\mathsf{T}}$ and assume $a(\lambda)_{\mathscr{r}}=0$. By construction,
@@
Step 2: By Step 1, every right zero $\lambda_{i}$ yields a right eigenpair $A_{c}v=v\lambda_{i}$. Hence the class $\lambda_{i}]$ belongs to the right spectrum. Indeed, if $A_{c}x=x\lambda_{i}$, then for every $q\in\mathbb{H}^{\times}$,
$A_{c}(xq)$ $=$ $(A_{c}x)q=x\lambda_{i}q=(xq)(q^{-1}\lambda_{i}q),$
so every element of $\lambda_{i}]$ is a right eigenvalue of $A_{c}$.
For the converse, take a right eigenpair $A_{c}x=x\mu$, $x\neq 0$, and write $x=[x_{1},\dots,x_{n}]^{\mathsf{T}}$. From the first n1n-1 rows,
$x_{k+1}$ $=$ $x_{k}\mu,\qquad k=1,\dots,n-1,$
hence
$x_{k}$ $=$ $x_{1}\mu^{k-1},\qquad k=1,\dots,n.$
Moreover, $x_{1}\neq 0$, since otherwise the recursion would give $x_{2}=\cdots=x_{n}=0$, contradicting $x\neq 0$.
Using the last row of $A_{c}x=x\mu$ and the above relations, we obtain
$-(a_{0}x_{1}+a_{1}x_{1}\mu+\cdots+a_{n-1}x_{1}\mu^{n-1})$ $=$ $x_{1}\mu^{n}.$
Right-multiplying by $x_{1}^{-1}$ gives
$-a_{0}-a_{1}(x_{1}\mu x_{1}^{-1})-\cdots-a_{n-1}(x_{1}\mu^{n-1}x_{1}^{-1})$ $=$ $x_{1}\mu^{n}x_{1}^{-1}.$
Let $\lambda:=x_{1}\mu x_{1}^{-1}$. Then $x_{1}\mu^{k}x_{1}^{-1}=\lambda^{k}$ for all $k$, so the equality becomes
$-a_{0}-a_{1}\lambda-\cdots-a_{n-1}\lambda^{n-1}=\lambda^{n}\;\iff\;a(\lambda)_{\mathscr{r}}=0.$
Thus every right eigenvalue $\mu$ is similar to a right zero $\lambda$ of $a(\lambda)$, so the class $\mu]$ equals $\lambda]$. Therefore every element of $\sigma_{\mathscr{r}}(A_{c})$ is of the form $\lambda]$ with $a(\lambda)_{\mathscr{r}}=0$. Together with the first inclusion, this yields
$\sigma_{\mathscr{r}}(A_{c})$ $=$ $\{[\lambda]:a(\lambda)_{\mathscr{r}}=0\}.$
Step 3: The left zeros of $a(\lambda)$ are known to be similar to its right zeros [8]. By Step 2, the right spectrum of $A_{c}$ contains the entire similarity class of every right zero. Hence every left zero of $a(\lambda)$ is a right eigenvalue of $A_{c}$ as well. ∎
@@
Throughout this note, we adopt the lower controllable companion form. Other orientations are related by permutation similarity, and therefore yield the same right spectrum. In this paper we fix the lower orientation, which uniquely determines the associated companion polynomial for a controllable pair.
@@
    Consequently,
    $\sigma_{\mathscr{r}}(A)=\{[\lambda]:a(\lambda)_{\mathscr{r}}=0\}.$
@@
Hence, the companion polynomial shares some properties of the minimal/characteristic polynomial even in quaternions, but not all! In particular, it does not annihilate $A$ in general; it only annihilates its companion form $A_{c}$. Hence $a(A_{c})_{\mathscr{r}}=0$ but, in general, $a(A)_{\mathscr{r}}\neq 0$.
@@
For a fixed controllable pair $(A,B)$, the associated lower companion realization is unique, and therefore so is its monic companion polynomial. When $A$ is cyclic, one may also speak informally of a companion polynomial of $A$, but this polynomial is not intrinsic to $A$ alone: different cyclic choices of $B$ generally produce different companion polynomials. What is invariant is the set of similarity classes determined by their right zeros, namely the right spectrum $\sigma_{\mathscr{r}}(A)$.